OpenStudy (anonymous):

what is the area of the triangle bounded by the lines: x=1, y=x-1, y=3-x

OpenStudy (anonymous):

Some thing doesn't seem right. Check and write question word for word. Why is there two y's?

OpenStudy (anonymous):

All the info is correctly listed. I'm still stumped on it.

OpenStudy (anonymous):

What is the area of the triangle bounded by the lines x = 1, y = x − 1, and y = 3 − x ?

OpenStudy (anonymous):

OK, I get it. These are graphs of lines, not actual measurements.

OpenStudy (anonymous):

Yes - they form a triangle when graphed. I need to find the area of that triangle.

OpenStudy (anonymous):

You can graph it and get their measurements. Or you can use algebra, for example \[x-1=3-x\]Solve for x. That is one point of intersection. Depends on if you are a visual person, draw graph, if you are a logical thinker and algebra lover, do it by algebra.

OpenStudy (anonymous):

so would the answer be 2? do I stop there?

OpenStudy (anonymous):

why just those 2 only?

OpenStudy (anonymous):

the 3rd one really bumped me off the train lol

OpenStudy (anonymous):

No that is just one point of the triangle, x=2, punch that in the other triangle y=1. One corner of the triangle is at the point (2,1)

OpenStudy (anonymous):

Alright, I worked it out and got (4) for the area of the triangle bounded by the lines

OpenStudy (anonymous):

were you able to work up a quick answer to compare?

OpenStudy (anonymous):

I haven't done it but if you tell me what you did, I can check.

OpenStudy (anonymous):

I found the other 2 corners of the triangle and added/subtracted them together and got (4).

OpenStudy (anonymous):

I don't know exactly what you did. You would have to give more detail. But to go that route you would have to know the (x,y) point of each intersection, use distance formula to compute distance then, area formula half base times height.

OpenStudy (anonymous):

I think I mixed up what I"m supposed to do

OpenStudy (anonymous):

I found all the points - where do I go from there exactly? I appreciate the help.

OpenStudy (anonymous):

This is a simple enough graph that you could plot by hand to get the distances.

OpenStudy (anonymous):

I just graphed it - just by looking at it, I see the area being a 1x1 box

OpenStudy (anonymous):

well 2 triangles

OpenStudy (anonymous):

so the area is = 1

OpenStudy (anonymous):

That is your first clue that you did something wrong. A triangle is a triangle, a box is a box. Take each y line, write down some values x=0, y= etc, and the line that says x=1 is just a straight line joining the two others.

OpenStudy (anonymous):

I did that - I cut the triangle in half and made a box and just counted the units

OpenStudy (anonymous):

My points for the triangle are: (1,1) - (2,1) - (1,2)

OpenStudy (anonymous):

don't mind the "-", I Just used those to space them out

OpenStudy (anonymous):

May be you are right, I am just not familiar with your method. Area of a triangle can be found by half base times height http://mathcentral.uregina.ca/QQ/database/QQ.09.02/dean1.html

OpenStudy (anonymous):

(1/2)(2) x (1) = 1

OpenStudy (anonymous):

Can we conclude that 1 is the answer?

OpenStudy (anonymous):

So what course are you working so I can understand what level of math you are or else I might say thing you might not understand.

OpenStudy (anonymous):

Calc

OpenStudy (anonymous):

Don't get hang up on any one question, write the number down and go in to your teacher's office hours. You have a good work ethic. Keep it up. But you have to put in a lot of work because you are not asking calculus questions. Don't get discouraged. 'Keep working,' that's what my tutor says at the end of each session.